OpenStudy (anonymous):

Domains {2,3,4,5} Ranges {4,7,10,13} write an equation for the relation given

OpenStudy (anonymous):

you have to use all the numbers?

OpenStudy (anonymous):

No, just and equation is needed.

OpenStudy (anonymous):

Okay I'm just guessing.. 2x+4y=8?

OpenStudy (anonymous):

Give me a minute to check that please

OpenStudy (anonymous):

Alrighty..

OpenStudy (anonymous):

What grade?

OpenStudy (anonymous):

I dont think so

OpenStudy (anonymous):

9th Algebra 1

OpenStudy (anonymous):

They need me to write an equation for the relation between X and Y

OpenStudy (anonymous):

I'm not sure what your doing.. what is it called? Like im in algebra 2 and we just got through with domains and range but we called it function stuff

OpenStudy (anonymous):

Yes same, They need me to write an equation for the relation between x and y

OpenStudy (anonymous):

Yes same, They need me to write an equation for the relation between x and y

OpenStudy (anonymous):

The relation between X an Y

OpenStudy (anonymous):

That could be anything though... Idk you got some hard stuff lol

OpenStudy (anonymous):

Yes same, They need me to write an equation for the relation between x and y

OpenStudy (anonymous):

Y= and ahead

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Yes same, They need me to write an equation for the relation between x and y

OpenStudy (anonymous):

A Y= equation. Whats the relation between X and Y

OpenStudy (anonymous):

it depends on what type of formula you want to use. which one of these looks most familiar to you as to what you use in class: y=mx+b y-y1=m(x-x1) ax + by = c

OpenStudy (anonymous):

Yes same, They need me to write an equation for the relation between x and y

OpenStudy (anonymous):

X 2 3 4 5 y 4 7 10 13

OpenStudy (anonymous):